What is bragg equation.


1
Expert's answer
2021-02-06T10:04:33-0500

Answer

The bragg equation is written as

"2d\sin\theta =n\lambda"

Where d is interplaner distance.

"\theta" is incident angle.

"\lambda" is wavelength of incident wave.